OpenStudy (calculusxy):

Steve, Jerry, and Ron were paid $29.25 to remove garden gnomes. They each worked four hours, except for Ron, who was 45 minutes late. How much of the $29.25 should Ron receive? @AllTehMaffs

OpenStudy (anonymous):

So if we set up the equation S+J+R=29.25 Can you define S, J, and R in terms of some variable x?

OpenStudy (calculusxy):

I don't understand what you mean by that?

OpenStudy (anonymous):

Like, x would be equivalent to an hourly work rate that we're trying to find, since we have the amount of time that all three of them worked. ie, S=4x because he worked for four hours, so would get paid 4hours *x dollars/hour

OpenStudy (calculusxy):

Is it like you multiply 4*2=8 and then you just do the adding of the 0.45. That would be equivalent to $8.45. I saw that that was the correct answer.

OpenStudy (anonymous):

I don't understand what you mean now ^^ Given this equation S+J+R=29.25 (does that part make sense?) I'm trying define those three terms in relation to some other variable, so instead of three unknowns in the equation, there's only be 1. So we can say S=4x J=4x R=3.25x (because 3 hours and 15 minutes is 3.25 hours)

OpenStudy (anonymous):

where "x" has units of dollars/hour

OpenStudy (anonymous):

Then the equation \[S+J+R=$29.25\] turns into \[(4 \ \text{hours})x+(4 \ \text{hours})x+(3.25 \ \text{hours})x=$29.25\] So you solve for x, which will be in units of $/h then to find out how much money each of them made, you just plug that x back into \[S=4x\] \[J=4x\] \[R=3.25x\]

OpenStudy (anonymous):

@calculusxy does that make sense, or have I lost you?